# Can RLHF be More Efficient with Imperfect Reward Models? A Policy Coverage Perspective

### v1

@@ Abstract
Sample efficiency is critical for online Reinforcement Learning from Human Feedback (RLHF). While existing works investigate sample-efficient online exploration strategies, the potential of utilizing misspecified yet relevant reward models to accelerate learning remains underexplored. This paper studies how to transfer knowledge from those imperfect reward models in online RLHF. We start by identifying a novel property of the KL-regularized RLHF objective: a policy’s ability to cover the optimal policy is captured by its sub-optimality. Building on this insight, we propose a theoretical transfer learning algorithm with provable benefits compared to standard online learning. Our approach achieves low regret in the early stage by quickly adapting to the best available source reward models without prior knowledge of their quality, and over time, it attains an $\widetilde{O}(\sqrt{T})$ regret bound independent of structural complexity measures. Inspired by our theoretical findings, we develop an empirical algorithm with improved computational efficiency, and demonstrate its effectiveness empirically in summarization tasks.
@@ 1 Introduction
    Firstly, policy value serves as a crucial criterion to select the transfer policy, since exploiting policies with high values does not conflict with exploration. Secondly, combining insights from offline RL theory, we prove that the offline policy, learned with online data generated by any no-regret online algorithm, converges to the optimal one at a rate of $\widetilde{O}(T^{-\frac{1}{2}})$ after a finite time. Such a bound improves existing sample complexity results and regret bounds in online RLHF by eliminating the dependencies on the size of the state and action space, or the complexity of the policy class. This suggests the offline policy computed with the data from the online learning process is a promising candidate to transfer from, which we term as “self-transfer learning”.
@@
  3. Inspired by the principles in TPO, we further propose an empirical algorithm in Sec. 5 with largely reduced computational overheads, making it scalable to practical RLHF scenarios. In particular, we employ an Upper-Confidence Bound (UCB) subroutine to select transfer policies with the highest win rates competing with the online learning policy. The effectiveness of our approach is demonstrated for fine-tuning T5 models on summarization tasks in Sec. 6.
@@ 1.1 Related Work
Figure 1: The standard pipeline for online RLHF (left) involves repeated online human feedback collection and policy improvement. Our setting additionally leverages imperfect reward models via transfer learning (right). Motivated by the the structure induced by KL regularization, we propose two principles for online RLHF transfer learning: (1) selecting transfer policy $\pi_{\text{Transfer}}$ with the highest policy value; (2) involving the offline policy $\pi_{\texttt{OFF}}$ computed with online collected data as a candidate (self-transfer learning).
@@ 2.1 Mathematical Formulation of RLHF
The policy class $\Pi$ satisfies: (I) Realizability: The optimal policy $\pi^{*}_{r^{*}}\in\Pi$. (II) Bounded Policy Ratio111 Note that $\max_{s,a}|\log\frac{\pi^{*}_{r^{*}}(a|s)}{\pi_{\texttt{ref}}(a|s)}|\leq\frac{R}{\beta}$ holds if $r^{*}\in[0,R]$ (Lem. B.1 in Appx. B.1). Thus, Assump. 2.1-(II) can be interpreted as an additional filtering step by leveraging the boundary of $r^{*}$.: $\forall\pi\in\Pi$, $\max_{s,a}|\log\frac{\pi(a|s)}{\pi_{\texttt{ref}}(a|s)}|\leq\frac{{R}}{\beta}.$
Additional Notation Given $\Pi$ satisfying Assump. 2.1, $\text{conv}(\Pi)$ denotes its convex hull, and $\mathcal{R}^{\Pi}$ denotes the reward class converted from $\Pi$, s.t., (1) $\forall r\in\mathcal{R}^{\Pi},r\in[0,R]$; (2) $\exists r\in\mathcal{R}^{\Pi},\pi_{r}^{*}=\pi^{*}_{r^{*}}$. We defer to Appx. B.1 for detailed converting process. Besides, we denote $[n]:=\{1,2,...,n\}$ and $a\wedge b:=\min\{a,b\}$. We refer the reader to Appx. A for a table of commonly used notation in this paper.
@@ 2.2 Reward Transfer Setting
LLM Policy as Reward Model  Eq. (2) implies that there is a one-to-one correspondence between policy and reward model. Given a LLM policy $\pi$ and an arbitrary distribution $\pi_{0}$, $\beta\log\frac{\pi(\cdot|\cdot)}{\pi_{0}(\cdot|\cdot)}$ can be interpreted as a reward model that $\pi$ aligns with given $\pi_{0}$ as the reference policy (Rosset et al., 2024). Therefore, while we consistently use the term “reward transfer” throughout the paper for clarity, our framework is general to handle transfer learning from any LLM policy through the underlying reward function it aligns with.
@@ 2.3 Background on Policy Coverability
Given a dataset $\mathcal{D}$ generated by a policy $\pi^{\mathcal{D}}$, running RPO with any $\mathcal{R}$ including $r^{*}$ and $\widetilde{\Pi}$ yields $\pi_{\texttt{OFF}}$, s.t., $\forall\pi\in\widetilde{\Pi},J_{\beta}(\pi)-J_{\beta}(\pi_{\texttt{OFF}})=\widetilde{O}(e^{2{R}}{\texttt{Cov}}^{\pi|\pi^{\mathcal{D}}}|\mathcal{D}|^{-\frac{1}{2}}).$
@@ 3.1 Structural Property Induced by Regularization
Under Assump. 2.1, for any policy $\pi\in\text{conv}(\Pi)$,
@@
The key insight of Lem. 3.1 is that: for any prospective policy candidates $\pi\in\text{conv}(\Pi)$, its coverability of $\pi^{*}_{r^{*}}$, i.e., ${\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi}$, is controlled by its policy value gap. Here we consider the convex hull in order to incorporate all possible uniform mixture policies induced by $\Pi$. Intuitively, ${\texttt{Cov}}^{\widetilde{\pi}|\pi}$ becomes extremely large or even unbounded if there is a significant distribution shift between $\pi$ and $\widetilde{\pi}$. However, in the presence of regularization ($\beta>0$), we should only consider policies with bounded policy ratio relative to $\pi_{\texttt{ref}}$ (see Assump. 2.1-(II)), and exclude those (near-)deterministic ones from our policy candidate class $\Pi$, because none of them can be (near-)optimal. In other words, regularization leverages prior knowledge from $\pi_{\texttt{ref}}$ and enables a free policy filtration step before learning begins, ensuring that the remaining policies exhibit a favorable structure (Lem. 3.1).
@@ 3.2 New Insights for Transfer Learning in RLHF
Inspired by Lem. 3.1, we identify two novel principles for transfer learning in the context of RLHF in the following, which we will further explore in later sections.
@@
Principle 2: Transfer from the Offline Policy—the “Self-Transfer Learning”  We first introduce a key result by combining Lem. 3.1 and offline RLHF result in Lem. 2.3.
@@
Under Assump. 2.1, given an online dataset $\mathcal{D}$ generated [^2] See Cond. C.1 for the definition of data generation process. by a policy sequence $\{\pi^{t}\}_{t=1}^{T}\in\text{conv}(\Pi)$, w.p. $1-\delta$, running RPO with $\text{conv}(\Pi)$, $\mathcal{R}^{\Pi}$ and $\mathcal{D}$ yields $\pi_{\texttt{OFF}}$:
$J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{OFF}})\leq$ (4)
$\widetilde{O}\Big{(}e^{2{R}}\cdot\Big{(}1+\kappa(e^{\frac{2{R}}{\beta}})\cdot\sum_{t=1}^{T}\frac{J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{t})}{\beta T}\Big{)}\sqrt{\frac{1}{T}}\Big{)}.$
To understand the significance of Thm. 3.2, consider the case when $\{\pi^{t}\}_{t\in[T]}$ are produced by a no-regret online learning algorithm, such as XPO in Lem. 2.4. As a result, the term $\sum_{t=1}^{T}\frac{J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{t})}{\beta T}$ in Eq. (4) diminishes to 0 as $T$ increases. This implies that the offline policy $\pi_{\texttt{OFF}}$ converges to $\pi^{*}_{r^{*}}$ at a rate of $O(T^{-\frac{1}{2}})$, which does not depend on $|\mathcal{S}|,|\mathcal{A}|$ or other complexity measures such as ${\texttt{Cov}}_{\infty}(\Pi)$ in Lem. 2.4. This result not only strictly improves the sample complexity bounds [^3] Beyond sample complexity, a regret bound improved to $\widetilde{O}(\sqrt{T})$ for online RLHF can be established. We defer it to Coro. 4.4 after presenting our main results. for existing online RLHF algorithms (Xiong et al., 2024; Xie et al., 2024; Cen et al., 2024; Zhang et al., 2024), but also reveals a fundamental difference from the pure reward maximization setting, where lower bounds depending on those structural complexity factors have been established (Auer et al., 2002; Dani et al., 2008). We defer detailed discussions to Appx. E.3.
More importantly, the faster rate of the convergence of $\pi_{\texttt{OFF}}$ to $\pi^{*}_{r^{*}}$ also indicates the potential of using $\pi_{\texttt{OFF}}$ as a candidate for policy transfer. We term this regime as “self-transfer learning”, and refer $\pi_{\texttt{OFF}}$ as the “self-transfer policy”. Notably, $\pi_{\texttt{OFF}}$ continuously improves and converges to $\pi^{*}_{r^{*}}$ as the dataset grows, while the source policies $\{\pi^{*}_{r^{w}}\}_{w=1}^{W}$ retain fixed non-zero value gaps due to the imperfections in reward models $\{r^{w}\}_{w=1}^{W}$. This reveals another benefit of self-transfer learning: it helps to avoid being restricted by suboptimal source reward models.
@@ 4 Provably Efficient Transfer Learning
We will use $\mathbb{E}_{\rho,\pi}[r]:=\mathbb{E}_{s\sim\rho,a\sim\pi}[r(s,a)]$ as a short note. In Line 2 of Alg. 2, $N(w;\mathcal{D}):=\sum_{i\leq|\mathcal{D}|}{\mathbb{I}}[\pi^{i}=\pi^{*}_{r^{w}}]$ denotes the number samples collected with $\pi^{*}_{r^{w}}$ in the dataset, following the convention that $1/N(\cdot,\cdot)=+\infty$ if $N(\cdot,\cdot)=0$. In Line 2, we leverage RPO (Liu et al., 2024) to compute the self-transfer policy $\pi_{\texttt{OFF}}$ and a reward function ${\widehat{r}}_{\texttt{OFF}}$. To save space, we defer the details of RPO to Appx. C.
@@
// Pessimistic estimation for $J_{\beta}(\pi_{\texttt{OFF}})-J_{\beta}(\pi_{\texttt{ref}})$.
$\pi_{\texttt{OFF}},{\widehat{r}}_{\texttt{OFF}}\leftarrow{\texttt{RPO}}(\text{conv}(\Pi),\mathcal{R}^{\Pi},\mathcal{D},\eta)$ with $\eta=c\cdot(1+e^{{R}})^{-2}\sqrt{\frac{1}{|\mathcal{D}|}\log\frac{|\Pi|T}{\delta}}.$
${\widehat{V}}(\pi_{\texttt{OFF}};\mathcal{D})\leftarrow\mathbb{E}_{\rho,\pi_{\texttt{OFF}}}[{\widehat{r}}_{\texttt{OFF}}]-\mathbb{E}_{\rho,\pi_{\texttt{ref}}}[{\widehat{r}}_{\texttt{OFF}}]-\beta\text{KL}(\pi_{\texttt{OFF}}\|\pi_{\texttt{ref}})+\frac{1}{\eta}L_{\mathcal{D}}({\widehat{r}}_{\texttt{OFF}})-\frac{1}{\eta}L_{\mathcal{D}}({\widehat{r}}_{\texttt{MLE}})-2ce^{2{R}}\sqrt{\frac{1}{|\mathcal{D}|}\log\frac{|\Pi|T}{\delta}}.$
return $\arg\max_{\pi\in\{\pi^{*}_{r^{w}}\}_{w\in[W]}\cup\{\pi_{\texttt{OFF}}\}}\widehat{V}(\pi;\mathcal{D})$. // Selecting Transfer Policy by Estimated Value
@@
Pessimistic Estimation for $J_{\beta}(\pi_{\texttt{OFF}})-J_{\beta}(\pi_{\texttt{ref}})$  The main challenge in estimating the value of $\pi_{\texttt{OFF}}$ is that, $\pi_{\texttt{OFF}}$ is not fixed but changing and improving. The previous optimistic strategy is not applicable here, since the coverage of $\pi_{\texttt{OFF}}$ by the dataset is unclear, making it difficult to quantify the uncertainty in estimation via count-based bonus term. Fortunately, given that $\pi_{\texttt{OFF}}$ is improving over time, it is more important when it surpasses all the other source policies. Therefore, it suffices to construct a tight lower bound for $J_{\beta}(\pi_{\texttt{OFF}})-J_{\beta}(\pi_{\texttt{ref}})$; see line 2. By leveraging $\widehat{r}_{\texttt{MLE}}$ and the optimality of $(\pi_{\texttt{OFF}},{\widehat{r}}_{\texttt{OFF}})$ for the RPO loss, we can show:
Lemma 4.2 (Value Est Error for $\pi_{\texttt{OFF}}$).
@@
$\leq\widehat{V}(\pi_{\texttt{OFF}};\mathcal{D})\leq J_{\beta}(\pi_{\texttt{OFF}})-J_{\beta}(\pi_{\texttt{ref}}).$ (6)
Here $\pi_{\text{mix}}^{\mathcal{D}}:=\frac{1}{|\mathcal{D}|}\sum_{i\leq|\mathcal{D}|}\pi^{i}$ denotes the mixture policy. In the LHS, the coefficient takes minimum over two factors ${\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi_{\text{mix}}^{\mathcal{D}}}$ and $\sqrt{\mathcal{C}(\Pi)}/\alpha$, resulting from two different ways to estimate the value gap of $\pi_{\texttt{OFF}}$. According to offline RLHF theory (see Lem. C.2), $\pi_{\texttt{OFF}}$ is competitive with any $\pi\in\text{conv}(\Pi)$ well-covered by the dataset distribution, or equivalently, $J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\texttt{OFF}})=J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi)+\widetilde{O}(|\mathcal{D}|^{-\frac{1}{2}}{\texttt{Cov}}^{\pi|\pi_{\text{mix}}^{\mathcal{D}}})$. By choosing $\pi=\pi^{*}_{r^{*}}$, we obtain the first bound with the factor ${\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi_{\text{mix}}^{D}}$. Next, considering $\pi=\frac{1}{\alpha kN}\sum_{i\leq k,j\leq\alpha N}\pi^{i,j}$, the uniform mixture of policies generated by $\texttt{Alg}_{\texttt{OL}}$ so far, leads to the second bound involving $\sqrt{\mathcal{C}(\Pi)}/\alpha$. This also explains why we still involve normal online learning in TPO—to provide another safeguard for the quality of the transfer policy.

### v2

@@ Abstract
Sample efficiency is critical for online Reinforcement Learning from Human Feedback (RLHF). While existing works investigate sample-efficient online exploration strategies, the potential of utilizing misspecified yet relevant reward models to accelerate learning remains underexplored. This paper studies how to transfer knowledge from those imperfect reward models in online RLHF. We start by identifying a novel property of the KL-regularized RLHF objective: a policy’s coverability of the optimal policy is captured by its sub-optimality. Building on this insight, we propose novel transfer learning principles and a theoretical algorithm with provable benefits compared to standard online learning. Our approach achieves low regret in the early stage by quickly adapting to the best available source reward models without prior knowledge of their quality, and over time, it attains an $\widetilde{O}(\sqrt{T})$ regret bound independent of structural complexity measures. Empirically, inspired by our theoretical findings, we develop a win-rate-based transfer policy selection method with improved computational efficiency. Moreover, our empirical transfer learning technique is modular and can be integrated with various policy optimization methods, such as DPO, IPO and XPO, to further enhance their performance. We validate the effectiveness of our method through experiments on summarization tasks.
@@ 1 Introduction
    Firstly, policy value serves as a crucial criterion to select the transfer policy, since exploiting policies with high values does not conflict with exploration. Secondly, combining insights from offline RL theory, we prove that the policy distilled by offline learning techniques from the online data generated by any no-regret online algorithm, converges to the optimal one at a rate of $\widetilde{O}(T^{-\frac{1}{2}})$ after a finite time. Such a bound improves existing sample complexity results and regret bounds in online RLHF by eliminating the dependencies on the size of the state and action space, or the complexity of the policy class (up to log-covering number). This suggests the offline policy computed with the data from the online learning process is a promising candidate to transfer from, which we term as “self-transfer learning”.
@@
  3. To reduce computational overheads and improve scalablity to practical RLHF scenarios, in Sec. 5, we further propose an empirical algorithm that selects transfer policy based on the win rates competing with the learning policy, rather than relying on the estimated policy values. On the one hand, win rates can be estimated much more efficiently than the policy values. On the other hand, as justified by our theory, win rates help to identify lower bounds for the coverability of the optimal policy. Notably, our empirical TPO is general: its core transfer learning technique can be modularized and combined with various reward-model-free policy optimization methods, such as DPO (Rafailov et al., 2024), IPO (Azar et al., 2024), XPO (Xie et al., 2024), to boost their performance. The effectiveness of our approach is demonstrated in fine-tuning T5 models on summarization tasks in Sec. 6.
@@ 1.1 Related Work
Figure 1: The standard online RLHF pipeline only involves learning from online human feedback (left). Our setting additionally leverages available imperfect reward models via transfer learning (right). Inspired by the structure induced by KL regularization, we propose novel principles for transfer learning in online RLHF: (1) selecting transfer policy $\pi_{\text{Transfer}}$ with the highest policy value; (2) self-transfer learning—involving as a candidate the policy $\pi_{\text{Dstl}}$ distilled from online collected data by offline learning techniques.
@@ 2.1 Mathematical Formulation of RLHF
The policy class $\Pi$ satisfies: (I) Realizability: The optimal policy $\pi^{*}_{r^{*}}\in\Pi$. (II) Bounded Policy Ratio: For any $\pi\in\Pi$, $\max_{s,a}|\log\frac{\pi(a|s)}{\pi_{\texttt{ref}}(a|s)}|\leq\frac{{R}}{\beta}.$
Note that $\max_{s,a}|\log\frac{\pi^{*}_{r}(a|s)}{\pi_{\texttt{ref}}(a|s)}|\leq\frac{R}{\beta}$ holds as long as $r\in[0,R]$ (see Lem. B.1 in Appx. B.1). Thus, Assump. 2.1-(II) is not actually an assumption, but can be interpreted as an additional filtering step by leveraging the boundary of $r^{*}$.
Additional Notation We use the standard big-O notations and use $\widetilde{(\cdot)}$ (e.g. $\widetilde{O}$) to suppress logarithmic terms, including log-covering numbers. Given $\Pi$ satisfying Assump. 2.1, $\text{conv}(\Pi)$ denotes its convex hull, and $\mathcal{R}^{\Pi}$ denotes the reward class induced by $\Pi$ via Eq. (2), s.t., (1) $\forall r\in\mathcal{R}^{\Pi},r\in[0,R]$; (2) $\exists r\in\mathcal{R}^{\Pi},\pi_{r}^{*}=\pi^{*}_{r^{*}}$. We defer to Appx. B.1 for detailed converting process. Besides, we denote $[n]:=\{1,2,...,n\}$ and $a\wedge b:=\min\{a,b\}$. We refer the reader to Appx. A for a table of commonly used notation in this paper.
@@ 2.2 Reward Transfer Setting
LLM Policy as Reward Model  Eq. (2) implies that there is a way to convert a given LLM policy $\pi$ to a reward model. Concretely, by choosing an arbitrary distribution $\pi_{0}$, we can interpret $\beta\log\frac{\pi(\cdot|\cdot)}{\pi_{0}(\cdot|\cdot)}$ as a reward function that $\pi$ aligns with given $\pi_{0}$ as the reference policy (Rosset et al., 2024). Therefore, while we consistently use the term “reward transfer” throughout the paper for clarity, our framework is general to handle transfer learning from any LLM policy through the underlying reward function it aligns with.
@@ 2.3 Background on Policy Coverability
Given a dataset $\mathcal{D}$ generated by a policy $\pi^{\mathcal{D}}$, running RPO with any $\mathcal{R}$ including $r^{*}$ and $\widetilde{\Pi}$ yields $\widehat{\pi}$, s.t., $\forall\pi\in\widetilde{\Pi},J_{\beta}(\pi)-J_{\beta}(\widehat{\pi})=\widetilde{O}(e^{2{R}}{\texttt{Cov}}^{\pi|\pi^{\mathcal{D}}}|\mathcal{D}|^{-\frac{1}{2}}).$
@@ 3.1 Structural Property Induced by Regularization
Under Assump. 2.1 and assume $r^{w}\in[0,R]$ for all $w\in[W]$, then, for any policy $\pi\in\text{conv}(\Pi)\cup\{\pi^{*}_{r^{w}}\}_{w=1}^{W}$,
@@
The key insight of Lem. 3.1 is that: for any prospective candidates of the optimal policy (i.e., $\text{conv}(\Pi)$ [^1] Here we consider the convex hull in order to incorporate all possible uniform mixture policies induced by $\Pi$.) or any transfer candidates (i.e., $\{\pi^{*}_{r^{w}}\}_{w=1}^{W}$), its coverability of $\pi^{*}_{r^{*}}$ is controlled by its policy value gap. Intuitively, ${\texttt{Cov}}^{\widetilde{\pi}|\pi}$ becomes extremely large or even unbounded if there is a significant distribution shift between $\pi$ and $\widetilde{\pi}$. However, in the presence of regularization ($\beta>0$), we should only consider policies with bounded policy ratio relative to $\pi_{\texttt{ref}}$ (see Assump. 2.1-(II)), and exclude those (near-)deterministic ones from our policy candidate class $\Pi$, because none of them can be (near-)optimal. In other words, regularization leverages prior knowledge from $\pi_{\texttt{ref}}$ and enables a free policy filtration step before learning begins, ensuring that the remaining policies exhibit a favorable structure (Lem. 3.1).
@@ 3.2 New Insights for Transfer Learning in RLHF
In the online RLHF, the primary goal of exploration is to discover high-reward regions, i.e., the states and actions covered by the optimal policy. Therefore, in our reward transfer setup, we propose to transfer from the policy with the best coverage of $\pi^{*}_{r^{*}}$. Inspired by Lem. 3.1, we identify two novel principles for transfer learning for RLHF objective, which we will further explore in later sections.
@@
Principle 2: Transfer from the Policy Distilled from Online Data—the “Self-Transfer Learning”  We first introduce a key result by combining Lem. 3.1 and offline RLHF result in Lem. 2.3.
@@
Under Assump. 2.1, w.p. $1-\delta$, given an online dataset $\mathcal{D}$ generated [^2] See Cond. C.1 for the definition of data generation process. by a policy series $\{\pi^{t}\}_{t=1}^{T}\in\text{conv}(\Pi)$, running RPO with $\text{conv}(\Pi)$ and $\mathcal{R}^{\Pi}$ on $\mathcal{D}$ yields a distilled policy $\pi_{\text{Dstl}}$, such that,
$J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\text{Dstl}})\leq$ (4)
$\widetilde{O}\Big{(}e^{2{R}}\Big{(}1+\kappa(e^{\frac{2{R}}{\beta}})\sum_{t=1}^{T}\frac{J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{t})}{\beta T}\Big{)}\sqrt{\frac{1}{T}}\Big{)}.$
To understand the significance of Thm. 3.2, consider the case when $\{\pi^{t}\}_{t\in[T]}$ are produced by a no-regret online learning algorithm, such as XPO in Lem. 2.4. As a result, the term $\sum_{t=1}^{T}\frac{J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi^{t})}{\beta T}$ in Eq. (4) diminishes to 0 as $T$ increases. This implies that the policy $\pi_{\text{Dstl}}$ distilled from online data by offline learning techniques converges to $\pi^{*}_{r^{*}}$ at a rate of $O(T^{-\frac{1}{2}})$, which does not depend on $|\mathcal{S}|,|\mathcal{A}|$ or other complexity measures such as ${\texttt{Cov}}_{\infty}(\Pi)$ in Lem. 2.4. This result not only strictly improves the sample complexity bounds [^3] Beyond sample complexity, a regret bound improved to $\widetilde{O}(\sqrt{T})$ for online RLHF can be established. We defer it to Coro. LABEL:coro:sqrtT_reg after presenting our main results. for existing online RLHF algorithms (Xiong et al., 2024; Xie et al., 2024; Cen et al., 2024; Zhang et al., 2024), but also reveals a fundamental difference from the pure reward maximization setting, where lower bounds depending on those structural complexity factors have been established (Auer et al., 2002; Dani et al., 2008). We defer detailed discussions to Appx. E.3.
More importantly, the faster rate of the convergence of $\pi_{\text{Dstl}}$ to $\pi^{*}_{r^{*}}$ also indicates the potential of using $\pi_{\text{Dstl}}$ as a candidate for policy transfer. We term this regime as “self-transfer learning”, and refer $\pi_{\text{Dstl}}$ as the “self-transfer policy”. Notably, $\pi_{\text{Dstl}}$ continuously improves and converges to $\pi^{*}_{r^{*}}$ as the dataset grows, while the source policies $\{\pi^{*}_{r^{w}}\}_{w=1}^{W}$ retain fixed non-zero value gaps due to the imperfections in reward models $\{r^{w}\}_{w=1}^{W}$. This reveals another benefit of self-transfer learning: it helps to avoid being restricted by suboptimal source reward models.
@@ 4 Provably Efficient Transfer Learning
We will use $\mathbb{E}_{\rho,\pi}[r]:=\mathbb{E}_{s\sim\rho,a\sim\pi}[r(s,a)]$ as a short note. In Line 2 of Alg. 2, $N(w;\mathcal{D}):=\sum_{i\leq|\mathcal{D}|}{\mathbb{I}}[\pi^{i}=\pi^{*}_{r^{w}}]$ denotes the number samples collected with $\pi^{*}_{r^{w}}$ in the dataset, following the convention that $1/N(\cdot,\cdot)=+\infty$ if $N(\cdot,\cdot)=0$. In Line 2, we leverage RPO (Liu et al., 2024) to compute the self-transfer policy $\pi_{\text{Dstl}}$ and a reward function ${\widehat{r}}_{\text{Dstl}}$. To save space, we defer the details of RPO to Appx. C.
@@
// Pessimistic estimation for $J_{\beta}(\pi_{\text{Dstl}})-J_{\beta}(\pi_{\texttt{ref}})$.
$\pi_{\text{Dstl}},{\widehat{r}}_{\text{Dstl}}\leftarrow{\texttt{RPO}}(\text{conv}(\Pi),\mathcal{R}^{\Pi},\mathcal{D},\eta)$ with $\eta=c\cdot(1+e^{{R}})^{-2}\sqrt{\frac{1}{|\mathcal{D}|}\log\frac{|\Pi|T}{\delta}}.$
${\widehat{V}}(\pi_{\text{Dstl}};\mathcal{D})\leftarrow\mathbb{E}_{\rho,\pi_{\text{Dstl}}}[{\widehat{r}}_{\text{Dstl}}]-\mathbb{E}_{\rho,\pi_{\texttt{ref}}}[{\widehat{r}}_{\text{Dstl}}]-\beta\text{KL}(\pi_{\text{Dstl}}\|\pi_{\texttt{ref}})+\frac{1}{\eta}L_{\mathcal{D}}({\widehat{r}}_{\text{Dstl}})-\frac{1}{\eta}L_{\mathcal{D}}({\widehat{r}}_{\texttt{MLE}})-2ce^{2{R}}\sqrt{\frac{1}{|\mathcal{D}|}\log\frac{|\Pi|T}{\delta}}.$
return $\arg\max_{\pi\in\{\pi^{*}_{r^{w}}\}_{w\in[W]}\cup\{\pi_{\text{Dstl}}\}}\widehat{V}(\pi;\mathcal{D})$. // Selecting Transfer Policy by Estimated Value
@@
Pessimistic Estimation for $J_{\beta}(\pi_{\text{Dstl}})-J_{\beta}(\pi_{\texttt{ref}})$  The main challenge in estimating the value of $\pi_{\text{Dstl}}$ is that, $\pi_{\text{Dstl}}$ is not fixed but changing and improving. The previous optimistic strategy is not applicable here, since the coverage of $\pi_{\text{Dstl}}$ by the dataset is unclear, making it difficult to quantify the uncertainty in estimation via count-based bonus term. Fortunately, given that $\pi_{\text{Dstl}}$ is improving over time, it is more important when it surpasses all the other source policies. Therefore, it suffices to construct a tight lower bound for $J_{\beta}(\pi_{\text{Dstl}})-J_{\beta}(\pi_{\texttt{ref}})$; see line 2. By leveraging $\widehat{r}_{\texttt{MLE}}$ and the optimality of $(\pi_{\text{Dstl}},{\widehat{r}}_{\text{Dstl}})$ for the RPO loss, we can show:
Lemma 4.2 (Value Est Error for $\pi_{\text{Dstl}}$).
@@
$\leq\widehat{V}(\pi_{\text{Dstl}};\mathcal{D})\leq J_{\beta}(\pi_{\text{Dstl}})-J_{\beta}(\pi_{\texttt{ref}}).$ (6)
Here $\pi_{\text{mix}}^{\mathcal{D}}:=\frac{1}{|\mathcal{D}|}\sum_{i\leq|\mathcal{D}|}\pi^{i}$ denotes the mixture policy. In the LHS, the coefficient takes minimum over two factors ${\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi_{\text{mix}}^{\mathcal{D}}}$ and $\sqrt{\mathcal{C}(\Pi)}/\alpha$, resulting from two different ways to estimate the value gap of $\pi_{\text{Dstl}}$. According to offline RLHF theory (see Lem. C.2), $\pi_{\text{Dstl}}$ is competitive with any $\pi\in\text{conv}(\Pi)$ well-covered by the dataset distribution, or equivalently, $J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi_{\text{Dstl}})=J_{\beta}(\pi^{*}_{r^{*}})-J_{\beta}(\pi)+\widetilde{O}(|\mathcal{D}|^{-\frac{1}{2}}{\texttt{Cov}}^{\pi|\pi_{\text{mix}}^{\mathcal{D}}})$. By choosing $\pi=\pi^{*}_{r^{*}}$, we obtain the first bound with the factor ${\texttt{Cov}}^{\pi^{*}_{r^{*}}|\pi_{\text{mix}}^{D}}$. Next, considering $\pi=\frac{1}{\alpha kN}\sum_{i\leq k,j\leq\alpha N}\pi^{i,j}$, the uniform mixture of policies generated by $\texttt{Alg}_{\texttt{OL}}$ so far, leads to the second bound involving $\sqrt{\mathcal{C}(\Pi)}/\alpha$. This also explains why we still involve normal online learning in TPO—to provide another safeguard for the quality of the transfer policy.